Exercise 10-21

Presented below is selected financial information for two divisions of Samberg Brewing.

Supply the missing information for the lettered items. (Round minimum rate of return and ROI to 1 decimal place, e.g. 10.5% and round average operating assets to 0 decimal places e.g. 58,971.)

Lager
Lite Lager
Contribution margin
$499,700
$300,800
Controllable margin
199,900
(c)
Average operating assets
$
(a)
$1,200,700
Minimum rate of return
% (b)
12
%
Return on investment
20
%
% (d)
Residual income
$100,400
$204,200
Answer #1

Calculate missing amount :

Lager Lite lager
Contribution margin 499700 300800
Controllable margin 199900 (1200700*12%+204200) = 348284
Average operating assets 199900*100/20 = 999500 1200700
Minimum rate of return (199900-100400)*100/999500 = 10% 12%
Return on investment 20% 348284*100/1200700 = 29%
Residual income 100400 204200